6) Nana just retired at the age of 62 and expects to live until she is 85 years old. She has $402,000 in her retirement savings account. She is somewhat conservative with her money and expects to earn 6 percent during her retirement years. How much can she withdraw from her retirement savings at the end of each month if she plans to spend her last penny on the morning of her death? b) Nana has some extra cash on hand and decided to give you $3,000 today. You have decided to save this money so that you can gift it to your grandchildren 50 years from now. How much additional money will you have to gift to your grandchildren if you can earn an average of 8.5 percent annually instead of just 8 percent annually on your savings? cNana wants to use the rest of her extra cash to buy herself a new sports car for $41,750, and the finance office at the dealership has quoted her an 8.6 percent APR loan compounded monthly for 48 months to buy the car. What is the effective interest rate on this loan?
Answer #1

(a) Let the monthly withdrawals be $ K, Withdrawal Tenure = 85 - 62 = 23 years or (23 x 12) = 276 months, Annual Interest Rate = 6 % per annum or (6/12) = 0.5 % per month

Current Savings = $ 402000

Therefore, 402000 = K x (1/0.005) x [1-{1/(1.005)^(276)}]

K = 402000 / 149.5109789 ~ $ 2688.76

(b) Current Amount = $ 3000, Tenure = 50 years , Available Interest Rates = 8% and 8.5 %

Additional Money Earned = 3000 x [(1.085)^(30) - (1.08)^(30)] = $ 4486.78

(c) Annual Percentage Rate = 8.6 % (APR)

Monthly Percentage Rate = (8.6 / 12) = 0.7167 %

Effective Annual Rate = [1 + Monthly Percentage Rate]^(12) - 1 = (1.007167)^(12) - 1 =0.08947 or 8.947 % ~ 8.95 %